ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 188, 2012

Alon, N., O. Angel, I. Benjamini and E. Lubetzky Sums and products along sparse graphs

353

Angel, O. See N. Alon

Avni, N., S. Lim and E. Nevo On commensurator growth  259

Balandraud, É. An addition theorem and maximal zero-sum free

sets in ℤ/pℤ 405

Barge, M., H. Bruin, L. Jones and L. Sadun Homological Pisot substitutions and exact regularity

281

Benjamini, I. See N. Alon

Braun, G. and R. Göbel Splitting kernels into small summands  221

Brozos-Vázquez, M., G. Calvaruso, E. García-Río and S. Gavino-Fernández Three-dimensional Lorentzian homogeneous Ricci solitons

385

Bruin, H. See M. Barge Calvaruso, G. See M. Brozos-Vázquez Dong, P. D., S. Shao and X. D. Ye Product recurrent properties,

disjointness and weak disjointness

463

Fernández-Unzueta, M. Extension of multilinear maps defined on subspaces

301

Flores, J., F. L. Hernández, E. M. Semenov and P. Tradacete Strictly singular and power-compact operators on Banach lattices

323

García-Río, E. See M. Brozos-Vázquez Gavino-Fernández, S. See M. Brozos-Vázquez Gesztesy, F., A. Gomilko, F. Sukochev and Y. Tomilov On a

question of A. E. Nussbaum on measurability of families of closed linear operators in a Hilbert space

195

Göbel R. See G. Braun Gomilko, A. See F. Gesztesy Goodrick, J. and M. C. Laskowski The Schröder–Bernstein prop-

erty for weakly minimal theories

91

Hernández, F. L. See J. Flores Jiang, Y., S. Mitra and H. Shiga Quasiconformal motions and

isomorphisms of continuous families of Möbius groups

177

Jones, L. See M. Barge Kashina, Y., S. Montgomery and S.-H. Ng On the trace of the

antipode and higher indicators

57

Kopecká, E., D. Reem and S. Reich Zone diagrams in compact subsets of uniformly convex normed spaces

1

Kumar, A. and S. K. Ray Weighted estimates for the k-plane transform of radial functions on Euclidean spaces

25

Laskowski, M. C. See J. Goodrick Leibman, A. A canonical form and the distribution of values of gen-

eralized polynomials

131

Lim, S. See N. Avni Lubetzky, E. See N. Alon Mitra, S. See Y. Jiang Mohammadi, A. and A. Salehi Golsefidy Simultaneous Diophan-

tine approximation in non-degenerate p-adic manifolds

231

Montgomery, S. See Y. Kashina Nevo, E. See N. Avni Ng, S.-H. See Y. Kashina Pergher, P. L. Q. An improvement of the five halves theorem of

J. Boardman

431

Ray, S. K. See A. Kumar Reem, D. See E. Kopecká Reich, S. See E. Kopecká Sadun, L. See M. Barge Salehi Golsefidy, A. See A. Mohammadi Semenov, E. M. See J. Flores Shao, S. See P. D. Dong

Shen, Ch.-Y. Algebraic methods in sum-product phenomena  123

Shiga, H. See Y. Jiang da Silva, V. R. T. On ℤ_{2}-graded identities of the super tensor prod-

uct of UT_{2}(F) by the Grassmann algebra

441

Sukochev, F. See F. Gesztesy Tomilov, Y. See F. Gesztesy Tradacete, P. See J. Flores Ye, X. D. See P. D. Dong Zapolsky, F. Reeb graph and quasi-states on the two-dimensional

torus 111